[figure: 9.7 cm]

QUANTUM ANISOTROPIC XY MODEL ON SPATIALLY ANISOTROPIC SQUARE LATTICE

O. Derzhko^{a}, T. Verkholyak^{a}, R. Schmidt^{b} and J. Richter^{b} aInstitute for Condensed Matter Physics, 1 Svientsitskii St., L’viv-11, 79011, Ukraine bInstitut für Theoretische Physik, Univrsität Magdeburg, P.O. Box 4120, D-39016 Magdeburg, Germany

We consider the s = \frac{1}{2} anisotropic XY model on a spatially anisotropic square lattice within the frames of several different approaches: the linear spin-wave theory, 2D Jordan-Wigner fermionization and exact diagonalization completed by finite-size scaling. Our aim is to study the zero temperature behaviour of the long-range order as the exchange interaction anisotropy varies interpolating between isotropic XY model and Ising model and the spatial anisotropy varies interpolating between spatially isotropic square lattice and a collection of noninteracting chains. The 2D Jordan-Wigner fermionization approach similarly to the coupled cluster method and the correlated basis function method detects a quantum phase transition driven by the exchange interaction anisotropy. The conclusions of our study for disappearance of the long-range order as interchain interaction decreases are quite similar to what have been recently found for a plane of interacting spin-\frac{1}{2} isotropic Heisenberg chains. O.D. thanks the STCU for the support (project #1673).

[figure: 13.4 cm]

Subject category :

6. Theory of Magnetism

Presentation mode : poster

Corresponding author : T. Verkholyak

Address for correspondence : Institute for Condensed Matter Physics, 1 Svientsitskii St., L’viv-11, 79011, Ukraine

Email address : werch@icmp.lviv.ua